OpenStudy (anonymous):

A set of equations is given below: Equation A: c = 2d + 1 Equation B: c = 3d + 5 Which of the following is a step that can be used to find the solution to the set of equations? 2d + 1 = 3d + 5 2d = 3d + 5 2d + 1 = 3d 2d + 5 = 3d + 1

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

i didn't get the question, could you explain more what you wanna solve ?

OpenStudy (anonymous):

A set of equations is given below: Equation A: c = 2d + 1 Equation B: c = 3d + 5 Which of the following is a step that can be used to find the solution to the set of equations? 2d + 1 = 3d + 5 2d = 3d + 5 2d + 1 = 3d 2d + 5 = 3d + 1

OpenStudy (triciaal):

another case of equating expressions that have the same value example 6 = 4 + 2 6 = 3 + 3 6 = 6 always so 4 + 2 = 3 + 3